OpenStudy (blackstreet23):

How do you solve for 'I' in these cases? 10 log (I/10^-12) = 120 and 10 log (I/10) = 0 Thanks a lot for your time!

OpenStudy (blackstreet23):

@freckles @pooja195 @Agl202

OpenStudy (blackstreet23):

@IrishBoy123

OpenStudy (irishboy123):

is it \[10 log (\dfrac{I}{10^{-12}}) = 120\]? and \(\log_{10}\)??

OpenStudy (blackstreet23):

10 log (I/10^-12)

OpenStudy (blackstreet23):

and yes to the first one

OpenStudy (blackstreet23):

@IrishBoy123

OpenStudy (blackstreet23):

@Michele_Laino

OpenStudy (michele_laino):

hint: we have these steps: \[\Large \begin{gathered} \log \left( {\frac{I}{{{{10}^{ - 12}}}}} \right) = \frac{{120}}{{10}} \hfill \\ \hfill \\ \log \left( {\frac{I}{{{{10}^{ - 12}}}}} \right) = 12 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, if we apply the definition of logarithm, we can write this: \[\huge \frac{I}{{{{10}^{ - 12}}}} = {10^{12}}\]

OpenStudy (michele_laino):

hint: if I multiply both sides by \(10^{-12}\), I get: \[\huge \frac{I}{{{{10}^{ - 12}}}} \times {10^{ - 12}} = {10^{12}} \times {10^{ - 12}}\] please continue